Grade 9Algebra

if 2^ x * 3 ^ x+4 = 7^ x then find the value of the x in the equation

Boddu Vijaya Bhasker Reddy
9 Years agoGrade 9

1 Answer

Arijit
9 Years ago
2x*3x+4=7x
2x*3x*34=7x
6x*34=7x
34=(7/6)^x
take log at both side and you will get ans x=(4log3)/(log(7/6))